(3x+y-4)dx+(x+y-2)dy=0


1
Expert's answer
2021-06-15T09:07:46-0400

Solution

Stationary point from the system  3x+y-4 = 0, x+y-2 = 0

Subtracting this equations 2x-2 = 0 => x=1, y=1

New variables X = x-1, Y = y-1

"\frac{dY}{dX}=-\frac{3x\ +y\ -4}{x\ +\ y\ -\ 2}=-\frac{3X+Y}{X+Y}=-\frac{Y/X+3}{Y/X+1}"

For z = Y/X or Y = z*X expression of the derivative is "\frac{dY}{dX}=\frac{dz}{dX}X+z"

So "\frac{dz}{dX}X+z=-\frac{z+3}{z+1}"   =>  "\frac{dz}{dX}X=-\frac{z^2+2z+3}{z+1}"

According to separable variable method  "\frac{\left(z+1\right)dz}{z^2+2z+3}=-\frac{dX}{X}"  and "\int\frac{\left(z+1\right)dz}{z^2+2z+3}=-\int\frac{dX}{X}"

After multiplication by 2: ln(z2+2z+3)+2ln|X|=lnC

(z2+2z+3)*X2=C => Y2+2X*Y+3X2 = C => (y-1)2+2(x-1)(y-1)+3(x-1)2 = C

Answer

(y-1)2+2(x-1)(y-1)+3(x-1)2 = C